OpenStudy (anonymous):

What is the center and radius of the following circle: x2 + (y – 6)² = 50

OpenStudy (jhannybean):

formula for circle \[\large (x-\color{green}h)^2+(y-\color{green}k)^2=r^2\] your equation: \[\large (x-\color{green}0)^2-(y-\color{green}6)^2=50\]

OpenStudy (jhannybean):

can you relate your equation to the formula for the circle? center= (h , k) Radius = ?

OpenStudy (anonymous):

Center : 0, 6 Radius: 7.071

OpenStudy (jhannybean):

Believe so.

OpenStudy (jhannybean):

Ty :)